OpenStudy (anonymous):

What is the general form of the equation of a circle with center at (a, b) and radius of length m? x2 + y2 − 2ax − 2by + (a2 + b2 − m2) = 0 x2 + y2 + 2ax + 2by + (a2 + b2 − m2) = 0 x2 + y2 − 2ax − 2by + (a + b − m2) = 0 x2 + y2 + 2ax + 2by + a2 + b2 = -m2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (michele_laino):

by definition, a circle whose center is C=(a,b) and whose radius is r, is given by the subsequent formula: \[{\left( {x - a} \right)^2} + {\left( {y - b} \right)^2} = {r^2}\]

OpenStudy (anonymous):

So would it be C

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

sorry, I don't think it is C